OpenStudy (anonymous):

Chase has been playing a game where he can create towns and help his empire expand. Each town he has allows him to create 1.13 times as many villagers. The game gave Chase 4 villagers to start with. Help Chase expand his empire by solving for how many villagers he can create with 17 towns. Then explain to Chase how to create an equation to predict the number of villagers for any number of towns. Show your work and use complete sentences.

OpenStudy (anonymous):

help please? i have the equation just need help solving

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (phi):

it sounds like a geometric series.

OpenStudy (anonymous):

i think its an=a1r^(n-1)

OpenStudy (phi):

you need numbers for a1, and r any ideas?

OpenStudy (anonymous):

a17=4r^(18-1) ?

OpenStudy (phi):

ok what is r ?

OpenStudy (phi):

hint 1.13 times as many villagers.

OpenStudy (anonymous):

a1=1.13

OpenStudy (anonymous):

1.13*4?

OpenStudy (anonymous):

let number of towns be tlet number of villagers be v Each town he has allows him to create 1.13 times as many villagers,implies t=1.13v then since he starts with 4 villagers we have 17=1.13(v+4)⟹v=171.13−4 in general t=1.13(v+4)⟹v=t1.13−4

OpenStudy (anonymous):

use the equation t=1.13(v+4), plug in for t and solve for v

OpenStudy (anonymous):

@topstryker

OpenStudy (anonymous):

did this help

OpenStudy (anonymous):

its hard for me to follow... can we try and fill in the geometric series formula?...

OpenStudy (anonymous):

Well the way I see it is he started with 4 villager so I would see the equation to be 1.13x times 17+4 would equal the amount of villagers he could have which would be 23.21 as for the equation it would be 1.13x times 17 plus four equals the number of villagers total

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would that answer the question you think? @lupita1995

OpenStudy (anonymous):

i think so

OpenStudy (phi):

*** 1.13*4*** that gives you the answer after one "step" Each town he has allows him to create 1.13 times as many villagers so he starts with 1 town with 4 people his next town has 1.13 * (1.13*4) = 1.13^2 * 4 or 4*1.13^2 for the next town we multiply by 1.13 to get the number of people in this town (not sure how we fractions of a person... which makes this a very goofy question) but ignoring how goofy it is, we get an= 4 * 1.13^(n-1) when n = 1 (the first town) we get 4* 1.13^(1-1) 1.13^0 is 1 (anything to the 0 power is 1), and we get 4 people.

OpenStudy (anonymous):

1.13(17+4) does that answer the question? =23.21? @phi

OpenStudy (phi):

no. read my post carefully

OpenStudy (anonymous):

where does the 17 towns come into?

OpenStudy (phi):

the geometric series is defined for any n (n is the number of towns) \[ a_n = 4 \cdot 1.13^{n-1} \]

OpenStudy (anonymous):

and n=17? @phi

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (phi):

** and n=17 *** yes, if you want to find the number of people he can create with 17 towns, use n=17 in the formula posted above.